An investor invested a total of $1,400 in two mutal funds. One fund earned a 7% profit and the other earned a 2% profit. If the investors profit was $58 how much was invested in each mutal fund?
fund that earned 7%?
fund that earned 2%?

Answer by mananth(16946)

Part I 7.00% per annum ------------- Amount invested =x
Part II 2.00% per annum ------------ Amount invested = y =
1400
Interest----- 58

Part I 7.00% per annum ---x
Part II 2.00% per annum ---y
Total investment
x + y= 1400 -------------1
Interest on both investments
7.00% x + 2.00% y= 58
Multiply by 100
7 x + 2 y= 5800.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -9800.00
Add this to (2)
0 x -5 y= -4000
divide by -5
y = 800
Part I 7.00% $ 600
Part II 2.00% $ 800

CHECK
600 --------- 7.00% ------- 42.00
800 ------------- 2.00% ------- 16.00
Total -------------------- 58.00

m.ananth@hotmail.ca